# The fundamental localization phases in quasiperiodic systems: A unified framework and exact results

### v2

@@ IV Exactly solvable models for all fundamental localization phases
In this section, we further demonstrate the application of the universal results. By implementing these results, we construct new exactly solvable models with different types of mobility edges (MEs), including the QP optical Raman lattice model that hosts all seven fundamental phases. In particular, by combining the first two theorems (Sec.III.1 and Sec.III.2), we construct a model that hosts three pure phases and four mixed phases. We achieve this by utilizing the theorems to construct parameters that govern the existence of MEs and critical states. In the absence of critical states, we obtain two pure phases and their mixed phases: namely, pure extended, pure localized, and the mixture of extended and localized states, where the first two phases do not exhibit MEs, while the last one does. By introducing critical states, we obtain the remaining pure phase and three distinct mixed phases. Finally, we apply the third theorem in Sec. III.3 to derive the conditions for exact solvability of the QP optical Raman lattice model and to explain the full phase diagram of the model.
In the following, we construct the exactly solvable models hosting all fundamental MEs in Sec. IV.1 and the QP optical Raman lattice model hosting all seven fundamental phases in Sec. IV.2.
IV.1 The model hosting all fundamental MEs
We begin by constructing exactly solvable models hosting all fundamental MEs by manipulating the generalized incommensurate zeros (GIZs) in matrix elements and the incommensurately distributed zeros (IDZs) in the shared component. In the presence of the QP modulation on the hopping coupling, as shown in Fig. 4(a), we can anticipate the emergence of critical orbitals. In addition, the spin-independent component $\sigma_{0}$ in the GIZs or the shared component breaks chiral symmetry, indicating the existence of MEs. Thus, by combining critical states with broken chiral symmetry, we can construct a model that hosts all fundamental MEs. The unified Hamiltonian is given by
@@ IV.1 The model hosting all fundamental MEs
This model shows how to manipulate the GIZs to introduce the critical orbitals into the system.
Figure 4: Generalized incommensurate zeros generated critical states. (a) The quasiperiodic spin-conserved hopping in this model gives rise to generalized incommensurate zeros (GIZs), which introduces critical orbitals in the system. (b) New exactly solvable model with generalized incommensurate zeros in matrix elements: Left panel: The model exhibits analytic mobility edges (MEs) at $E_{c}=\pm\lambda^{2}/\mu$, marked by the solid lines, which separate the extended states ($E<|\lambda^{2}/\mu|$) with fractal dimension ($\mathrm{FD}$) approaching 1, and the rigorous critical states ($E>|\lambda^{2}/\mu|$) with $\mathrm{FD}$ approaching a value between 0 and 1. Right panel: The corresponding $\mathrm{FD}$ for the dual model. The MEs $E_{c}=\pm\lambda^{2}/\mu$ (solid lines) now separate the localized states ($E<|\lambda^{2}/\mu|$) and the rigorous critical states ($E>|\lambda^{2}/\mu|$). (c) Exactly solvable model hosting all MEs: Upper panel: The energy-dependent effective potential resulting from the incommensurately distributed zeros in the shared component, which now acts as the IDZs on the effective dressed particle. Lower panel: The $\mathrm{FD}$ of the eigenstates shown as a function of $V_{B}/t$ and energy $E/t$, with $\lambda_{0}/t=1$ and $V_{0}/t=1.5$. The mobility edges are indicated by the solid lines. All systems have a size of $L=2586$.
We first introduce an exactly solvable model that hosts MEs separating critical and extended states, with the dual counterpart hosting MEs separating critical and localized states. By setting $t=V_{0}=V_{B}=0$ and $\lambda_{0}=\lambda_{1}=\lambda$, the hopping and on-site matrices become
$\Pi_{j}=\mu V_{j}^{\mathrm{od}}\Lambda_{-}+\lambda\sigma_{x},\quad M_{j}=\lambda\sigma_{x}.$ (28)
Fig. 4(b) left subfigure shows the model exhibits MEs separating critical and extended states at $E_{c}=\pm\lambda^{2}/\mu$. For energies $E<|\lambda^{2}/\mu|$, the system exhibits extended states, while for energies $E>|\lambda^{2}/\mu|$, critical states are observed. These results can be derived analytically by investigating the dual counterpart, whose coupling matrix is degenerate ($|\Pi_{n}|=0$) and is given by
@@
$t_{n}^{\mathrm{eff}}=\mu V_{n}^{\mathrm{od}},\quad V_{n}^{\mathrm{eff}}=\lambda^{2}(2+2V_{n}^{\mathrm{d}})/E.$ (30)
Neglecting the constant term in the potential, this effective model can be viewed as 1D extended Aubry-André (EAA) model Hatsugai and Kohmoto (1990); Han et al. (1994); Liu et al. (2015) without uniform hopping, where the system is in the localized (critical) phase when the QP on-site potential (QP hopping) dominates. Therefore, the system is in the critical phase for $\mu>\lambda^{2}/|E|$ and in the localized phase for $\mu<\lambda^{2}/|E|$. The numerical results shown in Fig. 4(b) right subfigure are consistent with this behavior. The MEs, $E_{c}=\pm\lambda^{2}/\mu$, separate the critical states for $|E|<\lambda^{2}/\mu$ from the localized states for $|E|>\lambda^{2}/\mu$. These results can also be derived by directly applying Avila’s global theory, yielding the Lyapunov exponent
@@
These results in the dual space provide an analytic characterization in the original space. Specifically, for $E<|\lambda^{2}/\mu|$, the eigenstate with energy $E$ in real space is extended, with a correlation length $\xi_{e}=\gamma^{-1}(E)$, as localized states are transformed into extended states under the dual transformation. For $E>|\lambda^{2}/\mu|$, the eigenstate with energy $E$ remains critical.
We further introduce a new exactly solvable model which supports all three types of MEs. In addition to the MEs marking the critical-to-localized transition and critical-to-extended transition, the model also exhibits traditional MEs that separate the extended and localized states. By setting $\mu=\lambda_{1}=0$, the model becomes
$\Pi_{j}=t\Lambda_{+},\quad M_{j}=(V_{0}+2V_{B}V_{j}^{\mathrm{d}})\Lambda_{-}+\lambda_{0}\sigma_{x}.$ (32)
As guided by the principles, the incommensurately distributed zeros (IDZs) appear at the shared components $\sigma_{0}$ and $\sigma_{z}$, indicating the emergence of critical orbitals. As shown in Fig. 4(c), when $V_{B}/t$ is large enough, the spectrum consists of critical and localized states. The MEs $E_{c}=\pm 2t$ separate the critical states for $|E|<2t$ from the localized states for $|E|>2t$. For intermediate values of $V_{B}/t$, extended orbitals appear in the spectrum.
The model can be reduced to an effective 1D spinless model with the effective hopping and on-site potential given by
$t_{j}^{\mathrm{eff}}=t,\quad V_{j}^{\mathrm{eff}}=\lambda_{0}^{2}/(E-V_{0}-2V_{B}V_{j}^{\mathrm{d}}).$ (33)
This energy-dependent on-site potential, shown in the upper panel of Fig. 4(c), is key to understanding the phase diagram. When $|E-V_{0}|\leq 2V_{B}$, the system exhibits an energy-dependent unbounded potential, which acts as the IDZs, leading to only critical and localized orbitals. This explains why, when $V_{B}/t$ is sufficiently large, the system exhibits only critical and localized states. In this regime, the analytic Lyapunov exponent is given by
$\gamma(E)=\ln|\frac{|E/t|+\sqrt{(E/t)^{2}-4}}{2}\Big{|}.$ (34)
This directly provides the mobility edges, $E_{c}=\pm 2t$, which separate the critical states ($|E|<2t$) from the localized states ($|E|>2t$). This regime demonstrates the resonant coupling mechanism for critical states. Specifically, in the spinful QP chain, when the on-site flipped coupling is zero ($\lambda_{0}\to 0$), the spin-up and spin-down chains decouple. The spin-down chain produces the localized orbitals with energies $E_{j}=V_{0}+2V_{B}\cos(2\pi\alpha j)$, while the spin-up chain produces extended orbitals. and the spin-up chain produces the extended orbitals. Resonantly hybridizing these orbitals causes the effective on-site potential to diverge, leading to IDZs that generate critical states.
When $|E-V_{0}|>2V_{B}$, the effective on-site potential is finite, leading to extended and localized states in this regime. The LE is given by
@@
with $\chi_{B}=2V_{B}/t$, $\chi_{E}=EV_{B}/t^{2}$ and $\chi_{0}=(E-V_{0})/t$. A finite $\gamma(E)$ leads to localized states with localization length $\xi_{l}=\gamma^{-1}(E)$, while $\gamma(E)=0$ corresponds to extended states.
Finally, by combining transition condition for different effective potential and the corresponding LEs from Eq. (34) and Eq. (35), we determine the analytic MEs that separate extended, localized and critical states, as illustrated in Fig. 4(b) and summarized in Table 1.
@@
IV.2 The model containing the seven fundamental localization phases
We further identify the seven fundamental phases realized in the 1D quasiperiodic optical Raman lattice model (Wang et al., 2020b, 2022b), as illustrated in Fig. 5(a). This model is constructed by manipulating the IDZs in the shared components and the chiral symmetry. Specifically, in the presence of chiral symmetry, three pure phases are obtained. Further breaking this symmetry introduces mobility edges and four additional coexisting phases. The Hamiltonian for the system is given by
@@ IV.2 The model containing the seven fundamental localization phases
where $t_{0}$ and $t_{\mathrm{so}}$ denote spin-conserved and spin-flipped hopping coefficients, $M_{z}$ is the strength of QP Zeeman potential, and $\eta$ is the chiral parameter that controls the ratio of spin-dependent to spin-independent QP potentials, governing the extent of chiral symmetry breaking.
@@
Fig. 5(b) illustrates the seven fundamental phases by varying the QP Zeeman potential $M_{z}$ and the chiral parameter $\eta$. These results are obtained by numerically diagonalizing the QP optical Raman lattice model in Eq.(36) with $t_{\mathrm{so}}=0.8t_{0}$. We first examine two limiting cases: exact chiral symmetry ($\eta=1$) and complete chiral symmetry breaking ($\eta=0$), followed by a discussion of the intermediate value of $\eta$. In the case of complete chiral symmetry breaking $\eta=0$, corresponding to a purely spin-independent QP potential, the system exhibits extended and localized phases in the weak and strong quasiperiodic potential regimes, respectively. In this case, the transition energies remain outside the eigenstate spectrum, preserving the pure phases. For the intermediate value of $\eta$, a coexisting phase (L+E) emerges, with MEs separating extended and localized states. This behavior can be understood as follows: in the absence of spin-flipped hopping ($t_{\mathrm{so}}=0$), the model in Eq.(36) describes two decoupled spin-up and spin-down QP chains with opposite energy spectra, where each chain hosts either purely localized or extended phases. Turning on $t_{\rm so}$ couples the two chains, opening gaps within each and flattening the dispersion of original band states. This coupling localizes part of states in the moderate potential regime, leading to the formation of MEs. On the other hand, in the chiral symmetry limit $\eta=1$, which corresponds to purely QP Zeeman potential, distinct phases appear in the weak, moderate, and strong field regimes, corresponding to pure extended, critical, and localized phases, respectively. These results are consistent with the predictions of Theorem I (Sec.III.1) and previous studies Wang et al. (2016, 2020b). As indicated by theorem I and II, coexisting phases involving critical states (C+E, L+C, L+E+C) necessitate introducing generalized incommensurate zeros in matrix elements and breaking chiral symmetry. This corresponds to the presence of both spin-dependent and spin-independent QP potentials. Fig. 5(b) indicates that these phases occur for $\eta_{c}<\eta<1$, with $\eta_{c}\approx 0.43$.
@@
The emergence of the seven fundamental phases can be analytically predicted by combining the universal results and Avila’s global theory. By applying Theorem III in Sec. III.3, the system is reduced to a spinless 1D QP chain with nearest neighbor hopping when $\det|\Pi_{j}|=0$, which occurs when $|t_{0}|=|t_{\mathrm{so}}|$. We set $t_{0}=t_{\mathrm{so}}$ without losing generality. The system exhibits chiral symmetry at $\eta=1$, where only pure critical and localized phases exist when $t_{\rm so}=t_{0}$, and all three pure phases are present when $t_{\rm so}\neq t_{0}$. Introducing a spin-independent quasiperiodic potential ($\eta\neq 1$) breaks chiral symmetry, leading to another exactly solvable point at $\eta=1/2$, where MEs emerge within the original critical and localized phases for $t_{\rm so}=t_{0}$, giving rise to the L+C phase. Further, various coexisting phases with combinations of the three types of quantum states emerge for $\eta<1$ when $t_{\mathrm{so}}\neq t_{0}$. Therefore, both numerical and analytical results confirm the existence of the seven fundamental localization phases, establishing this system as a universal quantum platform for exploration of localization physics.
@@
Figure 5: Seven fundamental phases of localization physics realized in 1D quasiperiodic (QP) optical Raman lattice. (a) Model Illustration: The hopping coupling matrix includes both spin-conserved and spin-flipped hopping terms, while the on-site matrix incorporates both spin-dependent quasiperiodic (QP) potential coupled to $\sigma_{z}$ and spin-independent QP potential coupled to $\sigma_{0}$. (b) Phase Diagram: The phase diagram, shown as a function of the QP Zeeman potential $M_{z}$ and the chiral parameter $\eta$, reveals seven distinct phases. These consist of three pure phases: extended (E), localized (L), and critical (C), as indicated in the diagram. In addition, there are three coexisting two-phase regions: (L+E) for the coexistence of localized and extended states, (L+C) for localized and critical states, and (C+E) for critical and extended states. Moreover, a coexisting three-phase region (L+E+C) appears, where all three phases coexist. The phase diagram is obtained with the parameter set $t_{\mathrm{so}}=0.8t_{0}$. (c) Exact Solvable Points: (c1) When $\eta=0$, the entire spectrum is localized. (c2) For $\eta=0.5$, the spectrum splits into localized and critical states, with the mobility edges marked by solid lines at $E_{c}=\pm 2t_{0}$. States with $|E|<2t_{0}$ are critical, while those with $|E|>2t_{0}$ are localized. (c3) At $\eta=1$, the system exhibits pure localized and critical phases, with the transition point marked by the dashed line at $M_{z}=4t_{0}$. The system is in the localized (critical) phase when $M_{z}>4t_{0}$ ($M_{z}<4t_{0}$). The system size used in these calculations is $L=2586$.

### v4

@@ IV Exactly solvable models for all fundamental localization phases
From the above subsection we can see that the universal results provide a powerful guidance to construct new exactly solvable models with nontrivial localization physics. In this section, we present the further in-depth study in developing the highly novel models with exact solutions, which enable a comprehensive and unified characterization of all the basic types of MEs and pure phases in quasiperiodic lattices. In particular, we propose two classes of exactly solvable models. The first is 1D spin-selective QP (SSQP) lattice model which is shown to host all basic types of MEs, and the second is the QP optical Raman lattice model which hosts all the seven fundamental localization phases. The seven fundamental phases are the three pure phases (extended, critical, and localized), three coexisting phases for any two of the three types of states, and one coexisting phase with all three types of states.
In the following, we first study the the exactly solvable models for all types of fundamental MEs. Then we show that the QP optical Raman lattice model can host all the seven fundamental localization phases.
IV.1 The spin-selective QP lattice model for all fundamental MEs
Figure 4: Generalized incommensurate zeros generated critical states. (a) The quasiperiodic (QP) spin-conserved hopping in this model gives rise to generalized incommensurate zeros (GIZs), which introduce critical orbitals. (b) The spin-selective QP (SSQP) lattice at the exactly solvable regime. Left panel: The model exhibits analytic mobility edges (MEs) at $E_{c}=\pm\lambda^{2}/\mu$, marked by the solid lines, which separate the extended states ($E<|\lambda^{2}/\mu|$) with fractal dimension $\mathrm{FD}\rightarrow 1$, and the rigorous critical states ($E>|\lambda^{2}/\mu|$) with $\mathrm{FD}$ approaching a value between 0 and 1. Right panel: The corresponding $\mathrm{FD}$ for the dual model. The MEs $E_{c}=\pm\lambda^{2}/\mu$ (solid lines) now separate the localized states ($E<|\lambda^{2}/\mu|$) and the critical states ($E>|\lambda^{2}/\mu|$). (c) The SSQP model at another exactly solvable regime that realizes all basic types of MEs. Upper panel: The energy-dependent effective potential resulted from the incommensurately distributed zeros in the shared component, which gives incommensurately distributed divergent potential. Lower panel: The $\mathrm{FD}$ of the eigenstates shown versus $V_{B}/t$ and $E/t$, with $\lambda_{0}/t=1$ and $V_{0}/t=1.5$. The MEs are indicated by the solid lines. All systems have a size of $L=2586$.
The SSQP lattice model for spin-$1/2$ fermions includes several ingredients. First, the QP hopping or QP onsite potential of the model is considered only for fermions at one spin state, while fermions at another spin state are coupled through the uniform spin-flip transitions. Second, the GIZs are introduced to the hopping matrix $\Pi_{j}$ or the shared component of $M_{j}$, leading to the emergence of the critical states. Third, the spin-independent coupling term is applied to break chiral symmetry, such that MEs emerge. The Hamiltonian is taken in a generic form
@@ IV.1 The spin-selective QP lattice model for all fundamental MEs
We first consider the exactly solvable regime that hosts MEs separating critical and extended states, with its dual counterpart hosting MEs separating critical and localized states. In particular, we take the condition $t=V_{0}=V_{B}=0$ and $\lambda_{0}=\lambda_{1}=\lambda$. The SSQP model [Eq. (29)] has QP hopping only for the spin-down fermions, whereas the spin-up fermions are coupled through uniform spin-flip tunneling and on-site transitions. In this case, the coupling matrices reduce to
$\Pi_{j}=\mu V_{j}^{\mathrm{od}}\Lambda_{-}+\lambda\sigma_{x},\ M_{j}=\lambda\sigma_{x}.$ (30)
Fig. 4(b) left subfigure shows the model exhibits MEs separating critical and extended states at $E_{c}=\pm\lambda^{2}/\mu$. For energies $E<|\lambda^{2}/\mu|$, the system exhibits extended states, while for energies $E>|\lambda^{2}/\mu|$, critical states are observed. These results can be derived analytically by investigating the dual counterpart, whose coupling matrix is degenerate $|\Pi_{n}|=0$ (Sec. III.2) and is given by
@@
$t_{n}^{\mathrm{eff}}=\mu V_{n}^{\mathrm{od}},\ V_{n}^{\mathrm{eff}}=\lambda^{2}(2+2V_{n}^{\mathrm{d}})/E.$ (32)
Neglecting the constant term in the potential, this effective model can be viewed as 1D extended Aubry-André (EAA) model Hatsugai and Kohmoto (1990); Han et al. (1994); Liu et al. (2015) without uniform hopping, where the system is in the localized (critical) phase when the QP potential (QP hopping) dominates (see Appendix A for details). Therefore, the system is in the critical phase for $\mu>\lambda^{2}/|E|$ and in the localized phase for $\mu<\lambda^{2}/|E|$. The numerical results shown in Fig. 4(b) right subfigure are consistent with this analytic study. The MEs, $E_{c}=\pm\lambda^{2}/\mu$, separate the critical states for $|E|<\lambda^{2}/\mu$ from the localized states for $|E|>\lambda^{2}/\mu$. These results can also be obtained by applying Avila’s global theory, yielding the LE
@@
With the above results in the dual space, we can easily obtain the analytic characterization in the original space. Specifically, for $E<|\lambda^{2}/\mu|$, the eigenstate with energy $E$ in real space is extended, with a correlation length $\xi_{e}=\gamma^{-1}(E)$, as localized states are transformed into extended states under the dual transformation. For $E>|\lambda^{2}/\mu|$, the eigenstate with energy $E$ remains critical.
We further consider another more exotic exactly solvable regime that realizes all three basic types of MEs, namely the MEs separating extended from localized states, extended from critical states, and critical from localized states. This corresponds to $\mu=\lambda_{1}=0$, for which the QP onsite potential is applied solely to the spin-down fermions in the SSQP model. The spin-up fermions have a uniform hopping, and are coupled to spin-down states through the onsite spin-flip transitions. In this regime the coupling matrices simplify to
$\Pi_{j}=t\Lambda_{+},\ M_{j}=(V_{0}+2V_{B}V_{j}^{\mathrm{d}})\Lambda_{-}+\lambda_{0}\sigma_{x}.$ (34)
The phase diagram in Fig. 4(c) reveals all three fundamental MEs resulting from the interplay between IDZs in shared components ($\sigma_{z},\sigma_{0}$) and the broken chiral symmetry. When $V_{B}/t$ is sufficiently large, the spectrum exhibits only critical and localized states, with MEs $E_{c}=\pm 2t$ separating the critical states ($|E|<2t$) and the localized states ($|E|>2t$). For intermediate $V_{B}/t$, extended orbitals also appear in the spectrum.
The underlying physics of the emergence of the three types of MEs can be achieved by by reducing the SSQP model in the current regime [Eq. (34)] to an effective 1D spinless model, where the effective energy-dependent on-site potential induced by IDZs in shared components plays a key role
$t_{j}^{\mathrm{eff}}=t,\ V_{j}^{\mathrm{eff}}=\lambda_{0}^{2}/(E-V_{0}-2V_{B}V_{j}^{\mathrm{d}}).$ (35)
When $|E-V_{0}|\leq 2V_{B}$, the system exhibits an incommensurately distributed divergent potential, leading to the critical states. This explains why, when $V_{B}/t$ is sufficiently large, the system exhibits only critical and localized states. In this regime, the analytic LE is
$\gamma(E)=\ln|\frac{|E/t|+\sqrt{(E/t)^{2}-4}}{2}\Big{|},$ (36)
with which the MEs reads $E_{c}=\pm 2t$. This result exemplifies a resonant coupling mechanism for critical states: when the onsite coupling is zero ($\lambda_{0}=0$), the two spin states decouple. The spin-down states produce a set of localized orbitals with energies $E_{j}=V_{0}+2V_{B}\cos(2\pi\alpha j)$, while the spin-up states produce extended orbitals. Resonantly hybridizing these orbitals leads to divergent effective on-site potential, generating critical states.
When $|E-V_{0}|>2V_{B}$, the effective potential is finite, giving extended and localized states. The LE can be analytically obtained as
@@
with $\chi_{B}=2V_{B}/t$, $\chi_{E}=EV_{B}/t^{2}$ and $\chi_{0}=(E-V_{0})/t$. A finite $\gamma(E)$ gives the localized states with localization length $\xi_{l}=\gamma^{-1}(E)$, and $\gamma(E)=0$ corresponds to the extended states.
Finally, by combining transition conditions for different effective potential and the corresponding LEs from Eq. (36) and Eq. (37), we determine the analytic MEs that separate extended, localized and critical states, as illustrated in Fig. 4(c) and summarized in Table 1.
@@
IV.2 The model with the seven fundamental localization phases
We now show a highly novel result that the seven fundamental localization phases can be realized in the 1D QP optical Raman lattice model (Wang et al., 2020b), as illustrated in Fig. 5(a). This model is constructed by manipulating the IDZs in the shared components and the chiral symmetry. Specifically, in the presence of chiral symmetry, three distinct pure phases are obtained. Breaking chiral symmetry introduces MEs and four additional coexisting phases. The Hamiltonian reads
@@ IV.2 The model with the seven fundamental localization phases
where $t_{0}$ ($t_{\mathrm{so}}$) denote spin-conserved (spin-flipped) hopping coefficient, $M_{z}$ is QP potential strength, and $\eta$ is the chiral parameter that controls the ratio of spin-dependent (Zeeman) to spin-independent (chemical) QP potentials, governing the extent of chiral symmetry breaking.
@@
Fig. 5(b) shows the phase diagram versus $M_{z}$ and $\eta$, which shows the presence of the seven fundamental phases. These results are first obtained by numerically diagonalizing the model in Eq.(38) with $t_{\mathrm{so}}=0.8t_{0}$. We examine two limiting cases: exact chiral symmetry ($\eta=1$) and complete chiral symmetry breaking ($\eta=0$), followed by a discussion of the intermediate value of $\eta$. In the case of complete chiral symmetry breaking $\eta=0$, corresponding to a purely spin-independent QP potential, the system exhibits extended and localized phases in the weak and strong quasiperiodic potential regimes, respectively. In this case, the transition energies remain outside the eigenstate spectrum, preserving the pure phases. For the intermediate value of $\eta$, a coexisting phase (L+E) emerges, with MEs separating extended and localized states. This result can be understood as follows: in the absence of spin-flipped hopping ($t_{\mathrm{so}}=0$), the model in Eq.(38) describes two decoupled spin-up and spin-down QP chains with opposite energy spectra, where each chain hosts either purely localized or extended phases. Turning on $t_{\rm so}$ couples the two chains, opening gaps within each and flattening the dispersion of original band states. This coupling localizes part of states in the moderate potential regime, leading to the formation of MEs. On the other hand, in the chiral symmetry limit $\eta=1$, which corresponds to purely QP Zeeman potential, distinct phases appear in the weak, moderate, and strong field regimes, corresponding to pure extended, critical, and localized phases, respectively. These results are consistent with the predictions of Theorem I (Sec.III.1) and previous studies Wang et al. (2016, 2020b). As indicated by theorem I and II, coexisting phases involving critical states (C+E, L+C, L+E+C) necessitate introducing generalized incommensurate zeros in matrix elements and breaking chiral symmetry. This corresponds to the presence of both spin-dependent and spin-independent QP potentials. Fig. 5(b) indicates that these phases occur for $\eta_{c}<\eta<1$, with $\eta_{c}\approx 0.43$.
@@
The emergence of the seven fundamental phases can be analytically predicted by combining the universal results and Avila’s global theory. By applying Theorem III in Sec. III.3, the system is reduced to a spinless 1D QP chain with nearest-neighbor hopping when $\det|\Pi_{j}|=0$, which occurs when $|t_{0}|=|t_{\mathrm{so}}|$. We set $t_{0}=t_{\mathrm{so}}$ without losing generality. The system exhibits chiral symmetry at $\eta=1$, where only pure critical and localized phases exist when $t_{\rm so}=t_{0}$, and all three pure phases are present when $t_{\rm so}\neq t_{0}$. Introducing a spin-independent quasiperiodic potential ($\eta\neq 1$) breaks chiral symmetry, leading to another exactly solvable point at $\eta=1/2$, where MEs emerge within the original critical and localized phases for $t_{\rm so}=t_{0}$, giving rise to the L+C phase. Further, various coexisting phases with combinations of the three types of quantum states emerge for $\eta<1$ when $t_{\mathrm{so}}\neq t_{0}$. Therefore, both numerical and analytical results confirm the existence of the seven fundamental localization phases, establishing this system as a universal quantum platform for exploration of localization physics.
@@
Figure 5: Seven fundamental phases realized in 1D quasiperiodic (QP) optical Raman lattice. (a) Model illustration: The hopping coupling matrix includes both spin-conserved and spin-flipped hopping terms, while the on-site matrix incorporates both spin-dependent quasiperiodic (QP) potential coupled to $\sigma_{z}$ and spin-independent QP potential coupled to $\sigma_{0}$. (b) Phase diagram: The phase diagram, shown as a function of the QP Zeeman potential $M_{z}$ and the chiral parameter $\eta$, reveals seven distinct phases. These consist of three pure phases: extended (E), localized (L), and critical (C), as indicated in the diagram. In addition, there are three coexisting two-phase regions: (L+E) for the coexistence of localized and extended states, (L+C) for localized and critical states, and (C+E) for critical and extended states. Moreover, a coexisting three-phase region (L+E+C) appears, where all three phases coexist. The phase diagram is obtained with the parameter set $t_{\mathrm{so}}=0.8t_{0}$. (c) Exact Solvable Points: (c1) When $\eta=0$, the entire spectrum is localized. (c2) For $\eta=0.5$, the spectrum splits into localized and critical states, with the mobility edges marked by solid lines at $E_{c}=\pm 2t_{0}$. States with $|E|<2t_{0}$ are critical, while those with $|E|>2t_{0}$ are localized. (c3) At $\eta=1$, the system exhibits pure localized and critical phases, with the transition point marked by the dashed line at $M_{z}=4t_{0}$. The system is in the localized (critical) phase when $M_{z}>4t_{0}$ ($M_{z}<4t_{0}$). The system size used in these calculations is $L=2586$.